1/3

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

The Banshee (An Approach)

Der Werwolf The Werewolf

El Hechicero

tradukita de Max Knight  tradukita de Alexander Gross  tradukita de Hans Eiseneck

One night, a banshee slunk away from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

Ein Werwolf eines Nachts entwich von Weib und Kind und sich begab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge mich!“ A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

Un hechicero, con esposa vino un día a la fosa de su maestro, y al profesor pidió: “¿Me cuentas? ¡Por fa-

vor!”

The village teacher climbed up

straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

Der Dorfschulmeister stieg hinauf auf seines Blechschilds Messing-

knauf und sprach zum Wolf, der seine

Pfoten geduldig kreuzte vor dem Toten:

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was

ready.

El profesor sin vacilar subió y comenzó a hablar. Desde allá de su letrero se dirigió al hechicero:

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

�Der Werwolf“ – sprach der gute

Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s

nennt, den Werwolf, - damit hat’s ein

End!“ “Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

“El hechiuno” le explica, “el hechidos” después indica, “el hechitres” dice al fin, “con eso basta ¡chiquitín!”

…  …  …  …

2/3

The banshee marveled at the cases and writhed with pleasure, making fa-

ces, but said: “You did not add, so far, the plural to the singular!”

Dem Werwolf schmeichelten die

Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“ “I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

El hechicero, conmovido por lo que capta su oído, al maestro ruega, encantado, que le adjunte un quebrado.

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts wus-

ste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

“A ‘Will-be-wolf?’ It’s just too long:

‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly

wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

Pero el profesor confiesa que nada sabe con certeza: “Hay muchos hechos. Hechi-

ceros únicamente hay enteros.”

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the te-

acher.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

The Werewolf knew better–his sons still

slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

El brujo se volvió muy triste: “Mi cara mitad, ¿no la viste?” Ya que no era sabio con gratitud se despidió.

…  …  …  …

3/3

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en hispana de Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .